OpenStudy (anonymous):

A 4-kg particle at rest at t=0 experiences a force of F=4+7t Newtons in a constant direction. What is the displacement of the particle at 4.2 s?

OpenStudy (anonymous):

a=F/m =(4+7*t)/4 x=Vin*t + (a*t^2)/2 Vin=0 because of at rest at t=0 x=((4+7*t)*(t^2))/8 then you can figure out what is x be from here.